# Error Estimation for Adaptive Mesh Refinement in Droplet Simulations

### v1

@@ 5 Conclusion
We presented a simple yet effective strategy to compute the error estimate for droplet simulations with front tracking method. The one-dimensional model used for the simulations is accurate for shear-induced droplets in a co-flowing environment. The mixed form equations naturally incorporate a smooth flux as a part of the solution, which allows us to use the flux-based error estimation. The comparison of error evolution for 85% glycerol droplets in quiescent air is presented for a non-refined and adaptively refined grid. This comparison shows that the adaptively refined grid provides improved stability and accuracy in capturing the droplet interface. The one-dimensional model has been previously used for conducting uncertainty quantification (UQ) analysis for droplet atomization in hybrid rocket combustion [26]. The presented error estimate can further enhance the accuracy of the UQ analysis.

### v2

@@ 5 Conclusion
We presented a simple yet effective strategy to compute the error estimate for droplet simulations with a front-tracking method. The one-dimensional model used for the simulations is accurate for shear-induced droplets in a co-flowing environment. The mixed form equations naturally incorporate a smooth flux as part of the solution, enabling flux-based error estimation. The comparison of error evolution for 85% glycerol droplets in co-flowing air is presented for unrefined, regularly refined, and adaptively refined grids. This comparison shows that the adaptively refined grid provides improved stability and accuracy in capturing the droplet interface. The AMR approach keeps the quantities of interest within $1\%$ of the regular-refinement case while reducing the computational cost by $\approx 76\%$ by using $\approx 82\%$ fewer elements.